# Loss Landscape Degeneracy Drives Stagewise Development in Transformers

### v2

@@ A.3.2 LLC estimation details for in-context linear regression
C Sampler # of chains $10$
$T_{\mathrm{SGLD}}$ Sampler # of SGLD steps / chain $5,000$
$\epsilon$ SGLD Step size $0.0003$
$\tilde{\gamma}=\epsilon\gamma/2$ SGLD Localization strength $0.01$
$\tilde{\beta}=\epsilon\beta/2n$ SGLD Inverse temperature $0.01$
(Default: $\beta^{*}=\frac{1}{\log n}$)
$m$ SGLD The size of each SGLD batch $2^{10}$
@@ C Sampler # of chains $10$
A.4 A guide to SGLD-based LLC estimation
@@
Figure A.1: Past some threshold, the choice of validation set size from which SGLD batches are sampled has little effect on learning coefficient estimates. Estimation hyperparameters are $C=8,T_{\mathrm{SGLD}}=2,000,m=2^{10},\epsilon=0.0003,\tilde{\gamma}=0.01,\tilde{\beta}=0.01$. Loss is evaluated over gradient minibatches at a representative selection of checkpoints. LLCs quickly converge to a constant value as the size increases.
@@ A.4.1 Varying the temperature
Since $\beta$ always shows up in a product with $n$ (in 8 for the SGLD step and in 3 for the LLC), we can view the inverse temperature as a multiplier that adjusts the effective size of your dataset. In a Bayesian setting, $\beta=2$ would mean updating twice on each of the samples in your dataset.
The problem with the default choice of $\beta^{*}$ is that as we increase $n$ we have to decrease the SGLD step size $\epsilon$ to prevent the update from becoming ill-conditioned, and this eventually causes the gradient term to suppress the noise term. This, in turn, leads to requiring larger batches to suppress the gradient noise and requiring longer chains to sufficiently explore the local posterior (Section A.4.3).
@@
A.4.2 Seeding the random noise
To smooth out the $\hat{\lambda}_{t}$ curves, we reset the random seed before LLC estimation run at each checkpoint. This means the sequence of injected Gaussian noise is the same for LLC estimation runs at different checkpoints. Additionally, if the batch size is held constant, the batch schedule will also be constant across different estimation runs. Figure A.3 shows that this does not affect the overall shape of the learning coefficient curves; it simply smooths it out.
@@ A.4.2 Seeding the random noise
A.4.3 Calibrating $\epsilon$, $\beta$, and $\gamma$
@@
where $\tilde{\beta}=\varepsilon\beta n/2$, $\tilde{\gamma}=\varepsilon\gamma/2$.
The results of this hyperparameter sweep are illustrated in Figure A.4 for final checkpoints. Separately (not pictured), we check the resulting hyperparameters for a subset of earlier checkpoints. This is needed since, for example, a well-behaved set of hyperparameters at the end of training may lead to failures like divergent estimates (Figure A.5) earlier in training when the geometry is more complex and thus the chains less stable.
Figure A.4: Results of grid sweep over SGLD hyperparameters for model 0 at $t=500\mathrm{k}$.
A.4.4 LLC traces
As a useful diagnostic when calibrating the LLC estimates, we propose an online variant for learning coefficient estimation. When overlaid on top of individual-chain LLC traces, this helps reveal common failure modes like divergent estimates, non-converged estimates, and escapes (Figure A.5). These traces display the running estimate of $\hat{\lambda}$ as a function of the number of steps taken in a chain (with the estimate averaged across independent chains).
@@ A.4.4 LLC traces
A.5 LLC estimates for a non-log-likelihood-based loss
@@
using the notation of Section 3 so $S_{k}=(x_{1},y_{1},\ldots,x_{k-1},y_{k-1},x_{k})$ is a context of length $k$. It is straightforward to check that this negative log likelihood $L$ agrees with the population loss $\ell$ associated to the empirical loss defined in Section 3. However the empirical quantities $L_{n}(w)$ and $\ell_{n}(w)$ defined for a set of samples of size $n$ are not the same.
Since we use the empirical loss $\ell_{n}$ in our calculation of the estimated LLC, whereas the foundational theory of SLT is written in terms of the empirical negative log likelihood $L_{n}$, it is natural to wonder how much of a difference this makes in practice. Figure A.6 depicts LLC traces (Section A.4) for a highlighted number of checkpoints using either a likelihood-based estimate (with variable sequence length) or loss-based estimate (with fixed sequence length). The relative orderings of complexities does not change, and even the values of the LLC estimates do not make much of a difference, except at the final checkpoint, which has a higher value for the sub-sequence-based estimate.
Figure A.6: Loss-based (left) and likelihood-based (right) LLC estimation yield identically ordered LLC estimates. With the exception of final checkpoint’s LLC estimate (which is larger for the loss-based estimate), the values are close to identical. These plots display LLC traces, which show the LLC estimate as a function of SGLD steps. This is a useful tool for calibrating LLC estimation (Section A.4).
A.6 LLC estimates away from local minima
@@ B.1 Procedure for stage boundary identification
To identify stage boundaries, we look for plateaus in the LLC: checkpoints at which the slope of $\hat{\lambda}(w_{t})$ over $t$ vanishes. To mitigate noise in the LLC estimates, we first fit a Gaussian process with some smoothing to the LLC-over-time curve. Then we numerically calculate the slope of this Gaussian process with respect to $\log t$. The logarithm corrects for the fact that the learning coefficient, like the loss, changes less as training progresses. We identify stage boundaries by looking for checkpoints at which this estimated slope equals zero. The results of this procedure are depicted in Figure B.1 for language and Figure B.2 for in-context linear regression.
@@ B.2 Stage boundary identification details for language model
Figure B.1 displays the test loss and LLC curves from Figure 1(a) in addition to the weight norm over time and associated slopes. Stage boundaries coincide with where the slope of the LLC crosses zero, that is, where there is a plateau in the LLC.
Figure B.1: A more detailed version of Figure 1(a) for two-layer language models. Top: Loss, LLC, and weight norm, along with an overlaid Gaussian process fit to these curves (red dotted lines). Bottom: Associated slopes, both numerically estimated finite differences (transparent blue) and of the Gaussian process (red dotted lined). Note that stage LM5 may be subdivided into further stages (Section B.1). However, the noise in LLC estimates late in training is high, so we do not draw any conclusions from this.
@@ B.3 Stage boundary identification details for in-context linear regression
Figure B.2 displays the test loss and LLC curves from Figure 1(b) in addition to the weight norm over time, and numerically estimated slopes associated to these three metrics. As in the case of language models, we identify stage boundaries by looking for plateaus in the LLC. Unlike the language models, here the boundaries LR1LR2 and LR2LR3 are clearly visible in the loss.
Figure B.2: A more detailed version of Figure 1(b) for in-context linear regression. Top: Loss, LLC, and weight norm, along with an overlaid Gaussian process fit to these curves (red dotted lines). Bottom: Associated slopes, both numerically estimated finite differences (transparent blue) and of the Gaussian process (red dotted lined). Top middle: Error bars displaying the standard deviation over the 10 SGLD chains are displayed in the background. Note that large error bars across chains are to be expected. Between different SGLD estimations, the variance is much lower. For example, averaged over training, the standard deviation over different seeds is only 4.2.
@@ B.4 Stage identification for additional training runs
Figure B.3: Figure 1(a) and Figure 1(b) for multiple seeds. In both settings, LLC reveals a consistent set of stages across five seeds. Late-training behavior shows more variance across seeds (see Sections B.4 and B.4).
Figure 3(a) shows loss and LLC curves for five seeds (differing in model initialization and batch schedule). In each seed, LLC estimation reveals stage LM1LM4. In three of the five seeds, stage LM5 is subdivided into two additional stages.
Figure 3(b) shows loss and LLC curves for five unique seeds (differing in model initialization and batch schedule). In each seed, LLC estimation reveals stages LR1LR5. There is remarkably little variance across different seeds.
@@ B.5 Comparison to Hessian statistics
Figure B.4 shows a quantification of the curvature-based notion of flatness captured by the Hessian (in contrast to the degeneracy-based notion of flatness captured by the LLC) for our in-context linear regression transformer. To estimate the trace and maximum eigenvalues shown in this figure, we use the PyHessian library (Yao et al., 2020) over a batch of $m=1024$ samples.
@@ Appendix C Developmental analysis of language models
In this section, we present further evidence on behavioral (Section C.1) and structural (Section C.2) development of the language model over the course of training.
@@ C.1.1 Bigram score
where we take fixed random sequences of $k_{i}$ and $S_{i}$ for $0001\leq i\leq n=5,$ 000, which is displayed over training in Figure 4(a). This is compared against the best-achievable bigram score, which is the bigram distribution entropy itself, averaged over the validation set.
@@ C.1.2 $n$-gram scores
In stage LM2 we consider $n$-grams, which are sequences of $n$ consecutive tokens, meaning $2$-grams and bigrams are the same. Specifically, we consider common $n$-grams, which is defined heuristically by comparing our 5,000 vocab size tokenizer with the full GPT-2 tokenizer. We use the GPT-2 tokenizer as our heuristic because its vocabulary is constructed iteratively by merging the most frequent pairs of tokens.
We first tokenize the tokens in the full GPT-2 vocabulary to get a list of 50,257 $n$-grams for various $n$. The first 5,000 such $n$-grams are all $1$-grams, after which $2$-grams begin appearing, then $3$-grams, $4$-grams, and so on (where $2$-grams and $3$-grams may still continue to appear later in the vocabulary). We then define the set of common $n$-grams as the first 1,000 $n$-grams that appear in this list for a fixed $n$, $n\geq 2$.
@@ C.1.3 In-context learning score
The in-context learning score is a behavioral measure of the relative performance of a model later in a sequence versus earlier in the sequence. We follow a similar construction as Olsson et al. (2022), where we take the loss at the 500th token minus the loss at the 50th token, so that a more negative score indicates better performance later in the sequence. This is then averaged over a 100k-row validation dataset. The performance of the language model over the course of training can be seen at the bottom of Figure 4(f).
@@ C.1.4 Visualizing behavioral changes
In Figure C.1, we visualize changes in the model’s input/output behavior by comparing model predictions before and after developmental stages and highlighting tokens with the greatest differences.
@@ C.2.1 Positional embedding
In Figure C.2, we measure the effect of the positional embedding on model performance by comparing the model’s performance at particular context positions on a validation set over the course of training against performance on the same validation set but with the positional embedding zero-ablated. The full context length is 1024, and we measure test loss at positions 1, 2, 3, 5, 10, 20, 30, 50, 100, 200, 300, 500, and 1000. In the transition from stage LM1 to LM2, the model begins using the learnable positional embedding to improve performance. The difference between test loss with and without the positional ablation is negligible at all measured positions until the LM1LM2 boundary.
@@
In Figure C.3, we examine the development of the positional embedding itself over time from two angles. The first is to take the embeddings of each position in the context and to run PCA on those embeddings. The result is that as training progresses, the positional embedding PCAs gradually resolve into Lissajous curves, suggesting that the positional embeddings might look like a random walk (Antognini and Sohl-Dickstein, 2018; Shinn, 2023). However, if we look to the explained variance, we see that it grows very large for PC1, reaching $94.2\%$ at training step 6400. This is much higher than we would expect for Brownian motion, where we expect to see about $61\%$ explained variance in PC1 (Antognini and Sohl-Dickstein, 2018).
@@ C.2.2 Composition scores
Where $M=W_{QK}^{h_{2}\ T}$, $M=W_{QK}^{h_{2}}$, and $M=W_{OV}^{h_{2}}$ for Q-, K-, and V-Composition respectively. See Figure C.4 for K-composition scores over time between attention heads in the induction circuits.
Figure C.4: The K-composition scores (Elhage et al., 2021) between first and second layer attention heads. The $h$th attention head in layer $l$ is indexed by $l:h$. The attention heads that eventually become previous token heads are $h=2,5$ in layer 1 (subplot rows 2 and 3), and the attention heads that eventually become induction heads are $h=7,8$ in layer 2 (subplot columns 2 and 3). The attention heads $1:1$ and $2:1$ are included for comparison. The induction heads $2:7$ and $2:8$ begin K-composing with first layer heads near the start of stage LM2. They continue to compose with the previous token heads in stages LM3 and LM4 (highlighted in green) while their K-composition scores drop with other attention heads in layer 1 in later stages.
@@ C.2.3 Previous-token matching score
The previous-token matching score is a structural measure of induction head attention. It is the attention score given to $[A]$ by an attention head at $[B]$ in the sequence $\ldots[A][B]$ (i.e., how much the head attends to the immediately preceding token).
@@
For each sample in this synthetic dataset, we measure the attention score that an attention head gives to the previous token when at the last token in the sequence. These scores are averaged across the dataset to produce the previous-token matching score for that attention head at a given checkpoint. The progression of previous-token matching scores over time can be seen in Figure 4(d).

### v3

@@ A.4.2 LLC estimation details for in-context linear regression
C Sampler # of chains 101010
$T_{\mathrm{SGLD}}$ Sampler # of SGLD steps / chain 500050005000
$\epsilon$ SGLD Step size 0.00030.00030.0003
$\gamma$ SGLD Localization strength 13.313.313.3
$n\beta$ SGLD Inverse temperature 66.766.766.7
$m$ SGLD The size of each SGLD batch 102410241024
@@ C Sampler # of chains 101010
A.5 A guide to SGLD-based LLC estimation
@@
Figure A.1: Past some threshold, the choice of validation set size from which SGLD batches are sampled has little effect on learning coefficient estimates. Estimation hyperparameters are $C=8,T_{\mathrm{SGLD}}=2,000,m=1024,\epsilon=0.0003,\tilde{\gamma}=0.01,\tilde{\beta}=0.01$. Loss is evaluated over gradient minibatches at a representative selection of checkpoints. LLCs quickly converge to a constant value as the size increases.
@@ A.5.1 Varying the temperature
Since $\beta$ always shows up in a product with $n$ (in ˜8 for the SGLD step and in ˜3 for the LLC), we can view the inverse temperature as a multiplier that adjusts the effective size of your dataset. In a Bayesian setting, $\beta=2$ would mean updating twice on each of the samples in your dataset.
The problem with the default choice of $\beta^{*}$ is that as we increase $n$ we have to decrease the SGLD step size $\epsilon$ to prevent the update from becoming ill-conditioned, and this eventually causes the gradient term to suppress the noise term. This, in turn, leads to requiring larger batches to suppress the gradient noise and requiring longer chains to sufficiently explore the local posterior (Section˜A.5.3).
@@
A.5.2 Seeding the random noise
To smooth out the $\hat{\lambda}_{t}$ curves, we reset the random seed before LLC estimation run at each checkpoint. This means the sequence of injected Gaussian noise is the same for LLC estimation runs at different checkpoints. Additionally, if the batch size is held constant, the batch schedule will also be constant across different estimation runs. Figure˜A.3 shows that this does not affect the overall shape of the learning coefficient curves; it simply smooths it out.
@@ A.5.2 Seeding the random noise
A.5.3 Calibrating $\epsilon$, $\beta$, and $\gamma$
@@
where $\tilde{\beta}=\varepsilon\beta n/2$, $\tilde{\gamma}=\varepsilon\gamma/4$.
The results of this hyperparameter sweep are illustrated in Figure˜A.4 for final checkpoints. Separately (not pictured), we check the resulting hyperparameters for a subset of earlier checkpoints. This is needed since, for example, a well-behaved set of hyperparameters at the end of training may lead to failures like divergent estimates (Figure˜A.5) earlier in training when the geometry is more complex and thus the chains less stable.
Figure A.4: Results of grid sweep over SGLD hyperparameters for model 0 at $t=27\mathrm{k}$.
A.5.4 LLC traces
As a useful diagnostic when calibrating the LLC estimates, we propose an online variant for learning coefficient estimation. When overlaid on top of individual-chain LLC traces, this helps reveal common failure modes like divergent estimates, non-converged estimates, and escapes (Figure˜A.5). These traces display the running estimate of $\hat{\lambda}$ as a function of the number of steps taken in a chain (with the estimate averaged across independent chains).
@@ A.5.4 LLC traces
A.6 LLC estimates for a non-log-likelihood-based loss
@@
using the notation of Section˜3 so $S_{k}=(x_{1},y_{1},\ldots,x_{k-1},y_{k-1},x_{k})$ is a context of length $k$. It is straightforward to check that this negative log likelihood $L$ agrees with the population loss $\ell$ associated to the empirical loss defined in Section˜3. However the empirical quantities $L_{n}(w)$ and $\ell_{n}(w)$ defined for a set of samples of size $n$ are not the same.
Since we use the empirical loss $\ell_{n}$ in our calculation of the estimated LLC, whereas the foundational theory of SLT is written in terms of the empirical negative log likelihood $L_{n}$, it is natural to wonder how much of a difference this makes in practice. Figure˜A.6 depicts LLC traces (Section˜A.5) for a highlighted number of checkpoints using either a likelihood-based estimate (with variable sequence length) or loss-based estimate (with fixed sequence length). The relative orderings of complexities does not change, and even the values of the LLC estimates do not make much of a difference, except at the final checkpoint, which has a higher value for the sub-sequence-based estimate.
Figure A.6: Loss-based (left) and likelihood-based (right) LLC estimation yield identically ordered LLC estimates. With the exception of final checkpoint’s LLC estimate (which is larger for the loss-based estimate), the values are close to identical. These plots display LLC traces, which show the LLC estimate as a function of SGLD steps. This is a useful tool for calibrating LLC estimation (Section˜A.5).
A.7 LLC estimates away from local minima
@@ B.1 Procedure for stage boundary identification
To identify stage boundaries, we look for plateaus in the LLC: checkpoints at which the slope of $\hat{\lambda}(w_{t})$ over $t$ vanishes. To mitigate noise in the LLC estimates, we first fit a Gaussian process with some smoothing to the LLC-over-time curve. Then we numerically calculate the slope of this Gaussian process with respect to $\log t$. The logarithm corrects for the fact that the learning coefficient, like the loss, changes less as training progresses. We identify stage boundaries by looking for checkpoints at which this estimated slope equals zero. The results of this procedure are depicted in Figure˜B.1 for language and Figure˜B.2 for in-context linear regression.
@@ B.2 Stage boundary identification details for language model
Figure˜B.1 displays the test loss and LLC curves from Figure˜1(a) in addition to the weight norm over time and associated slopes. Stage boundaries coincide with where the slope of the LLC crosses zero, that is, where there is a plateau in the LLC.
Figure B.1: A more detailed version of Figure˜1(a) for two-layer language models. Top: Loss, LLC, and weight norm, along with an overlaid Gaussian process fit to these curves (red dotted lines). Bottom: Associated slopes, both numerically estimated finite differences (transparent blue) and of the Gaussian process (red dotted lined). Note that stage LM5 may be subdivided into further stages (Section˜B.1). However, the noise in LLC estimates late in training is high, so we do not draw any conclusions from this.
@@ B.3 Stage boundary identification details for in-context linear regression
Figure˜B.2 displays the test loss and LLC curves from Figure˜1(b) in addition to the weight norm over time, and numerically estimated slopes associated to these three metrics. As in the case of language models, we identify stage boundaries by looking for plateaus in the LLC. Unlike the language models, here the boundaries LR1LR2 and LR2LR3 are clearly visible in the loss.
Figure B.2: A more detailed version of Figure˜1(b) for in-context linear regression. Top: Loss, LLC, and weight norm, along with an overlaid Gaussian process fit to these curves (red dotted lines). Bottom: Associated slopes, both numerically estimated finite differences (transparent blue) and of the Gaussian process (red dotted lined). Top middle: Error bars displaying the standard deviation over the 10 SGLD chains are displayed in the background. Note that large error bars across chains are to be expected. Between different SGLD estimations, the variance is much lower. For example, averaged over training, the standard deviation over different seeds is only 4.2.
@@ B.4 Stage identification for additional training runs
Figure B.3: Figure˜1(a) and Figure˜1(b) for multiple seeds. In both settings, LLC reveals a consistent set of stages across five seeds. Late-training behavior shows more variance across seeds (see Sections˜B.4 and B.4).
Figure˜3(a) shows loss and LLC curves for five seeds (differing in model initialization and batch schedule). In each seed, LLC estimation reveals stage LM1LM4. In three of the five seeds, stage LM5 is subdivided into two additional stages.
Figure˜3(b) shows loss and LLC curves for five unique seeds (differing in model initialization and batch schedule). In each seed, LLC estimation reveals stages LR1LR5. There is remarkably little variance across different seeds.
@@ B.5 Comparison to Hessian statistics
Figure˜B.4 shows a quantification of the curvature-based notion of flatness captured by the Hessian (in contrast to the degeneracy-based notion of flatness captured by the LLC) for our in-context linear regression transformer. To estimate the trace and maximum eigenvalues shown in this figure, we use the PyHessian library (Yao et al., 2020) over a batch of $m=1024$ samples.
@@ Appendix C Developmental analysis of language models
In this section, we present further evidence on behavioral (Section˜C.1) and structural (Section˜C.2) development of the language model over the course of training.
@@ C.1.1 Bigram score
where we take fixed random sequences of $k_{i}$ and $S_{i}$ for $1\leq i\leq n=5,000$, which is displayed over training in Figure˜4(a). This is compared against the best-achievable bigram score, which is the bigram distribution entropy itself, averaged over the validation set.
@@ C.1.2 $n$-gram scores
In stage LM2 we consider $n$-grams, which are sequences of $n$ consecutive tokens, meaning 222-grams and bigrams are the same. Specifically, we consider common $n$-grams, which is defined heuristically by comparing our 5,000 vocab size tokenizer with the full GPT-2 tokenizer. We use the GPT-2 tokenizer as our heuristic because its vocabulary is constructed iteratively by merging the most frequent pairs of tokens.
We first tokenize the tokens in the full GPT-2 vocabulary to get a list of 50,257 $n$-grams for various $n$. The first 5,000 such $n$-grams are all 111-grams, after which 222-grams begin appearing, then 333-grams, 444-grams, and so on (where 222-grams and 333-grams may still continue to appear later in the vocabulary). We then define the set of common $n$-grams as the first 1,000 $n$-grams that appear in this list for a fixed $n$, $n\geq 2$.
@@ C.1.3 In-context learning score
The in-context learning score is a behavioral measure of the relative performance of a model later in a sequence versus earlier in the sequence. We define $\text{ICL}_{k_{1}:k_{2}}$ to be the loss on token $k_{2}$ minus the loss on token $k_{1}$, so a more negative score indicates better relative performance later in the sequence. A more negative ICL score does not, however, mean that a model is achieving better overall loss on later tokens; it is only about the relative improvement. For the language model, we follow a similar construction as Olsson et al. (2022), where we take $k_{2}$ to be the 500th token and $k_{1}$ to be the 50th token. This is then averaged over a 100k-row validation dataset. The performance of the language model over the course of training can be seen at the bottom of Figure˜4(f).
@@ C.1.4 Visualizing behavioral changes
In Figure˜C.1, we visualize changes in the model’s input/output behavior by comparing model predictions before and after developmental stages and highlighting tokens with the greatest differences.
@@ C.2.1 Positional embedding
In Figure˜C.2, we measure the effect of the positional embedding on model performance by comparing the model’s performance at particular context positions on a validation set over the course of training against performance on the same validation set but with the positional embedding zero-ablated. The full context length is 1024, and we measure test loss at positions 1, 2, 3, 5, 10, 20, 30, 50, 100, 200, 300, 500, and 1000. In the transition from stage LM1 to LM2, the model begins using the learnable positional embedding to improve performance. The difference between test loss with and without the positional ablation is negligible at all measured positions until the LM1LM2 boundary.
@@
[…]
@@ C.2.2 Composition scores
Where $M=W_{QK}^{h_{2}\ T}$, $M=W_{QK}^{h_{2}}$, and $M=W_{OV}^{h_{2}}$ for Q-, K-, and V-Composition respectively. See Figure˜C.4 for K-composition scores over time between attention heads in the induction circuits.
Figure C.4: The K-composition scores (Elhage et al., 2021) between first and second layer attention heads. The $h$th attention head in layer $l$ is indexed by $l:h$. The attention heads that eventually become previous token heads are $h=2,5$ in layer 1 (subplot rows 2 and 3), and the attention heads that eventually become induction heads are $h=7,8$ in layer 2 (subplot columns 2 and 3). The attention heads $:1$ 1 and 2:12:12 : 1 are included for comparison. The induction heads 2:72:72 : 7 and 2:82:82 : 8 begin K-composing with first layer heads near the start of stage LM2. They continue to compose with the previous token heads in stages LM3 and LM4 (highlighted in green) while their K-composition scores drop with other attention heads in layer 1 in later stages.
@@ C.2.3 Previous-token matching score
[…]
@@
For each sample in this synthetic dataset, we measure the attention score that an attention head gives to the previous token when at the last token in the sequence. These scores are averaged across the dataset to produce the previous-token matching score for that attention head at a given checkpoint. The progression of previous-token matching scores over time can be seen in Figure˜4(d).